Rahdi played on online trivia quiz game. In the game , each right answer is worth 3 points, and each incorrect answer was worth -5 points. Randi answered 25 quiz questions and had a total of 27 points. How many more correct answers than incorrect answers did Randi have? I know the answer is 19 correct and 6 incorrect, but I am having trouble finding the correct equation.
Answer by stanbon(75887)

Let number of correct answers be "x"
The number of wrong answers is "25-x"
#of points from right answers is 3x
#of points from wrong answers is -5(25-x)=125-5x
EQUATION:
Pts + Pts= 27
3x-125+5x=27
8x=152
x=19 (# of correct answers)
25-x=6 (# of wrong answers)
Cheers,
Stan H.